# SAFREE: Training-Free and Adaptive Guard for Safe Text-to-Image And Video Generation

### v1

@@ 3.4 SAFREE for Advanced T2I Models and Text-to-Video Generation
Unlike existing unlearning-based methods limited to specific models or tasks, SAFREE is architecture agnostic and can be integrated across diverse backbone models without model modifications, offering superior versatility in safe generation. This flexibility is enabled by concept-orthogonal, selective token projection and self-validating adaptive filtering, allowing SAFREE to work across a wide range of generative models and tasks. It operates seamlessly with models beyond SD v-1.4, like, SDXL (Podell et al., 2023) and SD-v3 (stabilityai, 2024) in a zero-shot, training-free manner, and extends its applicability to text-to-video (T2V) generation models like ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024d), making it highly flexible as a series of plug-and-play modules. We present both qualitative and quantitative results showing its effectiveness across various model backbones (UNet (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023)) and tasks (T2I and T2V generation) in Sec. 4.6.
@@ 4.1 Experimental Setup
We use StableDiffusion-v1.4 (SD-v1.4) (Rombach et al., 2022) as the primary T2I backbone, following recent work (Gandikota et al., 2023; 2024; Gong et al., 2024). All methods are tested on adversarial prompts from red-teaming methods: I2P (Schramowski et al., 2023), P4D (Chin et al., 2024), Ring-a-Bell (Tsai et al., 2024), MMA-Diffusion (Yang et al., 2024b), and UnlearnDiff (Zhang et al., 2023b). Following Gandikota et al. (2023), we also evaluate models on artist-style removal tasks, using two datasets: one with five famous artists (Van Gogh, Picasso, Rembrandt, Warhol, Caravaggio) and the other with five modern artists (McKernan, Kinkade, Edlin, Eng, Ajin: Demi-Human), whose styles can be mimicked by SD. We extend SAFREE to text-to-video generation, applying it to ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024d) with different model backbones (UNet (Ronneberger et al., 2015) and Diffusion Transformer (Peebles & Xie, 2023)). For quantitative evaluation, we use SafeSora (Dai et al., 2024) with 600 toxic prompts across 12 concepts, constructing a benchmark of 296 examples across 5 categories.
@@
{tabu}
l—cccc—cccc Remove “Van Gogh” Remove “Kelly McKernan”
[…]
SD-v1.4–0.950.95–0.800.83
CA0.300.130.650.900.220.170.500.76
RECE0.310.080.800.930.290.040.550.76
UCE0.250.050.950.980.250.030.800.81
SLD-Medium0.210.100.950.910.220.180.500.79
SAFREE (Ours)0.420.310.350.850.400.390.400.78
4.2 Baselines and Evaluation Metrics
4.3 Evaluating the Effectiveness of SAFREE
SAFREE achieves training-free SoTA performance without altering model weights. We compare different methods for safe T2I generation, extensively and comprehensively evaluating each model’s vulnerability to adversarial attacks (i.e., attack success rate (ASR)) and their performance across multiple attack scenarios. As shown in Sec. 4.1 and Fig. 4, the proposed safeguard SAFREE consistently achieves significantly lower ASR than all training-free baselines, across all attack types. Notably, it demonstrates 47%, 13%, and 34% lower ASR compared to the best-performing counterparts— I2P, MMA-diffusion, and UnlearnDiff, respectively—highlighting its strong resilience against adversarial attacks.
Figure 4: Generated examples of SAFREE and safe T2I baselines. Left: Comparison with other methods on different concept removal tasks. Right: SAFREE incorporates with different T2I backbone models (SDXL and SD-v3) and T2V models (ZeroScopeT2V and CogVideoX). More visualizations are in the Appendix.
SAFREE show competitive results against training-based methods. In addition, we compare our approach with training-based methods. Surprisingly, our approach achieves competitive performance against these techniques. While SA and MACE exhibit strong safeguarding capabilities, they significantly degrade the overall quality of image generation due to excessive modifications of SD weights, often making them impractical for real-world applications as they frequently cause severe distortions. Notably, SAFREE delivers comparable safeguarding performance while generating high-quality images on the COCO-30k dataset, all within a training-free framework.
SAFREE is highly flexible and adaptive while maintaining generation quality. Our SAFREE does not require additional training or model weight modifications (more detailed comparison in later Tab. 4), providing key advantages over other methods (e.g., ESD, SA, and CA) which depend on unlearning or stochastic optimization, thereby increasing complexity. SAFREE allows for dynamic control of the number of filtered denoising steps based on inputs in an adaptable manner without extensive retraining or model modifications. Furthermore, it is worth noting that, compared with other methods, SAFREE is able to preserve the safe content within the original prompt through targeted filtering in both the token and pixel spaces, ensuring that the projected embeddings remain within the input space, which makes SAFREE a highly efficient and reliable solution for real-world applications. As shown in Fig. 4 left-top, when requiring removing the artist concept (“Van Gogh”), our SAFREE deletes this targeted art style (first row) while preserving other artists style (second row) by producing semantically similar outputs to the original SD. The examples of removing the “nudity” concept as shown in Fig. 4 left-bottom draw a similar conclusion. It demonstrates our SAFREE is a highly adaptive safe generation solution that is able to maintain untargeted safe concepts well.
Ablations. We validate the effectiveness of three components of SAFREE in Tab. 4 using SD-v1.4. First, we examine the impact of the adaptive toxic token selection (T, Sec. 3.1). We replace the selected token embeddings with either the null token embedding (N) or our proposed projected embeddings (P, Sec. 3.2). Both variants significantly reduce ASR from adversarial prompts, demonstrating the effectiveness of our toxic token selection. However, we observe that using N results in degraded image quality, as inserting null tokens disrupts the prompt structure and shifts the input embeddings outside their original space. In contrast, our orthogonal projection (P) achieves competitive safeguard performance with better COCO evaluation results. Incorporating self-validating filtering (S, Eq. 6) further enhances image quality by amplifying the filtering when input tokens are relevant to toxic concepts, although it can slightly reduce filtering capability. By integrating these components with latent-level re-attention (L, Sec. 3.3), our final model strikes a strong balance between effective, safe filtering and preserving the quality of images generated from prompts unrelated to toxic concepts.
Table 3: Ablation Study on the proposed SAFREE. T: Token Projection. S: Self-validating filtering. L: Latent Re-attention. N: Replacing with Null embedding. P: Orthogonal Token Projection.
@@ 4.3 Evaluating the Effectiveness of SAFREE
Table 5: Ours with SDXL and SD-v3. We include more visualization in Appendix.
@@
Table 6: SAFREE for safe video generation on SafeSora benchmark. Our method is applied on ZeroScopeT2V and CogVideoX-5B. We also include more visualization in Appendix.
@@
4.4 Evaluating SAFREE on Artist Concept Removal Tasks
As shown in Sec. 4.1, the proposed SAFREE achieves higher LPIPSe and LPIPSu scores compared to the baselines. The latter is likely due to our approach performing denoising processes guided by a coherent yet projected conditional embedding within the input space. As illustrated in Fig. 5, SAFREE enables generation models to retain the artistic styles of other artists very clearly even with larger feature distance (i.e., high LPIPSu). To validate whether the generated art styles are accurately removed or preserved, we frame these tasks as a multiple-choice QA problem, moving beyond feature-level distance assessments. Here, Acce and Accu represent the average accuracy of erased and unerased artist styles predicted by GPT-4o based on corresponding text prompts. As shown in  Sec. 4.1, SAFREE demonstrates impressive ability in removing targeted artist concepts, while baseline methods struggle to erase essential representations of target artists.
4.5 Efficiency of SAFREE
This section compares the efficiency of various methods, including the training-based ESD/CA, which update models through online optimization and loss, and the training-free UCE/RECE, which modify model attention weights using closed-form edits. Our method (SAFREE), similar to SLD, is training-free and filtering-based, without altering diffusion model weights. As shown in Tab. 4, while UCE/RECE offer fast model editing, they still require additional time for model updates. In contrast, SAFREE requires no model editing or modification, providing flexibility for model development across different conditions while maintaining competitive generation speeds. Based on Sec. 4.1 and Tab. 4, SAFREE delivers the best overall performance in concept safeguarding, generation quality, and flexibility.
@@ 4.6 Generalization and Extensibility of SAFREE
To further validate the robustness and generalization of SAFREE, we apply our method to various Text-to-Image (T2I) backbone models and Text-to-Video (T2V) applications. We extend SAFREE from SD-v1.4 to more advanced models, including SDXL, a scaled UNet-based model, and SD-V3, a Diffusion Transformer (Peebles & Xie, 2023) model. Our SAFREE demonstrates strong, training-free filtering of unsafe concepts, seamlessly integrating with these backbones. As shown in Tab. 6, SAFREE reduces unsafe outputs by 48% and 47% across benchmarks/datasets for SD-XL and SD-V3, respectively. We also extend SAFREE to T2V generation, testing it on ZeroScopeT2V (zeroscope, 2024) (UNet based) and CogVideoX-5B (Yang et al., 2024d) (Diffusion Transformer based) using the SafeSora (Dai et al., 2024) benchmark. As listed in Tab. 6, SAFREE significantly reduces a range of unsafe concepts across both models. It highlights SAFREE’s strong generalization across architectures and applications, offering an efficient safeguard for generative AI. This is also evident in Fig. 4 right, demonstrating that SAFREE with recent powerful T2I/T2V generation models can produce safe yet faithful (e.g., preserve the concept of ‘woman’ in CogVideoX + SAFREE) and quality visual outputs. More visualizations for T2I and T2V models are included in the Appendix.
@@ Acknowledgement
We thank Jaemin Cho, Zhongjie Mi, and Chumeng Liang for the useful discussion and feedback. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, and ONR Grant N00014-23-1-2356. The views contained in this article are those of the authors and not of the funding agency.
@@
  • Extra visualization of our methods and other baselines, T2I generation with other backbone models (SDXL and SD-v3), video generation results with T2V backbones (ZeroScopeT2V and CogVideoX) (Sec. B).
  • Limitations and Broader Impact of our proposed SAFREE (Sec. C).
  • License information (Sec. D).
Appendix A Experimental Details
@@ A.1 Experimental Setup
We employ StableDiffusion-v1.4 (SD-v1.4) (Rombach et al., 2022) as the main text-to-image generation backbone, following the recent literature (Gandikota et al., 2023; 2024; Gong et al., 2024). We evaluate our approach and baselines on inappropriate or adversarial prompts from multiple red-teaming techniques:
I2P (Schramowski et al., 2023): contains $4703$ unsafe prompts related to seven inappropriate/toxic concepts, including hate, harassment, violence, self-harm, sexual, shocking, and illegal activity. We evaluate our model and baselines by removing only the nudity concept following Gong et al. (2024), as the other concepts are ambiguous to determine their degree.
P4D (Chin et al., 2024): we use the P4D-N-16 dataset which contains $151$ unsafe prompts collected by a red-teaming tool, Prompting4Debugging (P4D). This is obtained by the prompt engineering to generate inappropriate concepts/objects from (safe) SD variants, including SD with negative prompts, ESD (Gandikota et al., 2023), and SLD (Schramowski et al., 2023).
Ring-a-bell (Tsai et al., 2024): we use $79$ problematic prompts collected by a model agnostic red-teaming framework, Ring-a-bell, which is built upon a text encoder (e.g., text encoder in CLIP model) and is executed offline without dependency on T2I models.
MMA-Diffusion (Yang et al., 2024b): is a red-teaming framework that is designed to generate inappropriate attack prompts by considering both textual and visual inputs to bypass defensive approaches for T2I models. We use $1$K strong adversarial prompts obtained by this framework.
UnlearnDiffAtk (Zhang et al., 2023b): we also use $142$ adversarial prompts generated by an adversarial prompt attack framework, named UnLearnDiffAtk for unlearned T2I diffusion models (i.e., diffusion models with additional unlearning on unsafe concepts/objects).
In addition to evaluating safe T2I generation, we further assess the models’ reliability in artist-style removal tasks. Following Gandikota et al. (2023), we employ two datasets: The first includes five famous artists: Van Gogh, Pablo Picasso, Rembrandt, Andy Warhol, and Caravaggio, while the second contains five modern artists: Kellly McKernan, Thomas Kinkade, Tyler Edlin, Kilian Eng, and Ajin: Demi-Human, whose styles have been confirmed to be imitable by SD. Each dataset contains $20$ prompts per artist style; therefore, it has $100$ prompts in total.
We further extend our SAFREE to text-to-video generation. We apply our method to two video generation models, ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024d) with different model backbones (UNet and Diffusion Transformer (Peebles & Xie, 2023)). To quantitatively evaluate the unsafe concept filtering ability on T2V, we choose SafeSora (Dai et al., 2024), which contains 600 toxic textual prompts across 12 toxic concepts as our testbed. We further select 5 representative categories within 12 concepts, and thus construct a safe video generation benchmark with 296 examples.

### v2

@@ 3.5 SAFREE for Advanced T2I Models and Text-to-Video Generation
Unlike existing unlearning-based methods limited to specific models or tasks, SAFREE is architecture agnostic and can be integrated across diverse backbone models without model modifications, offering superior versatility in safe generation. This flexibility is enabled by concept-orthogonal, selective token projection and self-validating adaptive filtering, allowing SAFREE to work across a wide range of generative models and tasks. It operates seamlessly with models beyond SD v-1.4, like, SDXL (Podell et al., 2023) and SD-v3 (stabilityai, 2024) in a zero-shot, training-free manner, and extends its applicability to text-to-video (T2V) generation models like ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024c), making it highly flexible as a series of plug-and-play modules. We present both qualitative and quantitative results showing its effectiveness across various model backbones (UNet (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023)) and tasks (T2I and T2V generation) in Sec. 4.6.
@@ 4.1 Experimental Setup
We use StableDiffusion-v1.4 (SD-v1.4) (Rombach et al., 2022) as the primary T2I backbone, following recent work (Gandikota et al., 2023; 2024; Gong et al., 2024). All methods are tested on adversarial prompts from red-teaming methods: I2P (Schramowski et al., 2023), P4D (Chin et al., 2024), Ring-a-Bell (Tsai et al., 2024), MMA-Diffusion (Yang et al., 2024a), and UnlearnDiff (Zhang et al., 2023b). Following Gandikota et al. (2023), we also evaluate models on artist-style removal tasks, using two datasets: one with five famous artists (Van Gogh, Picasso, Rembrandt, Warhol, Caravaggio) and the other with five modern artists (McKernan, Kinkade, Edlin, Eng, Ajin: Demi-Human), whose styles can be mimicked by SD. We extend SAFREE to text-to-video generation, applying it to ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024c) with different model backbones (UNet (Ronneberger et al., 2015) and Diffusion Transformer (Peebles & Xie, 2023)). For quantitative evaluation, we use SafeSora (Dai et al., 2024) with 600 toxic prompts across 12 concepts, constructing a benchmark of 296 examples across 5 categories.
4.2 Baselines and Evaluation Metrics
@@
Table 2: Ablation Study on SAFREE. T: Token Projection. S: Self-validating filtering. L: Latent Re-attention. N: Replacing with Null embedding. P: Orthogonal Token Projection.
@@
4.3 Evaluating the Effectiveness of SAFREE
SAFREE achieves training-free SoTA performance without altering model weights. We compare different methods for safe T2I generation, extensively and comprehensively evaluating each model’s vulnerability to adversarial attacks (i.e., attack success rate (ASR)) and their performance across multiple attack scenarios. As shown in Sec. 4.2 and Fig. 3, SAFREE consistently achieves significantly lower ASR than all training-free baselines across all attack types. Notably, it demonstrates 47%, 13%, and 34% lower ASR compared to the best-performing counterparts, I2P, MMA-diffusion, and UnlearnDiff, respectively, highlighting its strong resilience against adversarial attacks.
SAFREE shows competitive results against training-based methods. In addition, we compare our approach with training-based methods. Surprisingly, our approach achieves competitive performance against these techniques. While SA and MACE exhibit strong safeguarding capabilities, they significantly degrade the overall quality of image generation due to excessive modifications of SD weights, often making them impractical for real-world applications as they frequently cause severe distortions. Notably, SAFREE delivers comparable safeguarding performance while generating high-quality images on the COCO-30k dataset, all within a training-free framework.
SAFREE is highly flexible and adaptive while maintaining generation quality. SAFREE does not require additional training or model weight modifications (more detailed comparison in later Tab. 3), providing key advantages over other methods (e.g., ESD, SA, and CA) which depend on unlearning or stochastic optimization, thereby increasing complexity. SAFREE allows for dynamic control of the number of filtered denoising steps based on inputs in an adaptable manner without extensive retraining or model modifications. Furthermore, it is worth noting that, compared with other methods, compared to other methods, SAFREE preserves safe content in the original prompt through targeted joint filtering and ensures projected embeddings stay within the input space, making it highly efficient and reliable for real-world applications.
As shown in Fig. 3 left-top, when requiring removing the artist concept (“Van Gogh”), SAFREE deletes this targeted art style (first row) while preserving other artists style (second row) by producing semantically similar outputs to the original SD. The examples of removing the “nudity” concept as shown in Fig. 3 left-bottom draw a similar conclusion. It demonstrates SAFREE is a highly adaptive safe generation solution that is able to maintain untargeted safe concepts well.
Figure 3: Generated examples of SAFREE and safe T2I baselines. Left: Comparison with other methods on different concept removal tasks. Right: SAFREE incorporates with different T2I and T2V models. We provide more visualizations in the appendix (Sec. C).
Ablations. We validate the effectiveness of three components of SAFREE in Tab. 3 using SD-v1.4. First, we examine the impact of the adaptive toxic token selection (T, Sec. 3.1). We replace the selected token embeddings with either the null token embedding (N) or our proposed projected embeddings (P, Sec. 3.2). Both variants significantly reduce ASR from adversarial prompts, demonstrating the effectiveness of our toxic token selection. However, we observe that using N results in degraded image quality, as inserting null tokens disrupts the prompt structure and shifts the input embeddings outside their original space. In contrast, our orthogonal projection (P) achieves competitive safeguard performance with better COCO evaluation results. Incorporating self-validating filtering (S, Eq. 6) further enhances image quality by amplifying the filtering when input tokens are relevant to toxic concepts, although it can slightly reduce filtering capability. By integrating these components with latent-level re-attention (L, Sec. 3.4), our method strikes a strong balance between effective, safe filtering while preserving image quality for prompts unrelated to toxic concepts.
4.4 Evaluating SAFREE on Artist Concept Removal Tasks
As shown in Sec. 4.5, SAFREE achieves higher LPIPSe and LPIPSu scores compared to the baselines, where LPIPSe and LPIPSu denote the average LPIPS for images generated in the target erased artist styles and others (unerased), respectively. The higher LPIPSu score is likely due to our approach performing denoising processes guided by a coherent yet projected conditional embedding within the input space. As shown in Fig. 5, SAFREE enables generation models to retain the artistic styles of other artists very clearly even with larger feature distances (i.e., high LPIPSu). To validate whether the generated art styles are accurately removed or preserved, we frame these tasks as a multiple-choice QA problem, moving beyond feature-level distance assessments. Here, Acce and Accu represent the average accuracy of erased and unerased artist styles predicted by GPT-4o based on corresponding text prompts. As shown in  Sec. 4.5, SAFREE effectively remove targeted artist concepts, while baselines struggle to erase key representations of target artists.
4.5 Efficiency of SAFREE
We compare the efficiency of various methods, including the training-based ESD/CA, which update models through online optimization and loss, and the training-free UCE/RECE, which modify model attention weights using closed-form edits. Similar to SLD, our method (SAFREE) is training-free and filtering-based, without altering diffusion model weights. As shown in Tab. 3, while UCE/RECE offer fast model editing, they still require additional time for updates. In contrast, SAFREE requires no model editing or modification, providing flexibility for model development across different conditions while maintaining competitive generation speeds. Based on Sec. 4.2 and Tab. 3, SAFREE delivers the best overall performance in concept safeguarding, generation quality, and flexibility.
{tabu}
l—cccc—cccc Remove ”Van Gogh” Remove ”Kelly McKernan”
[…]
SD-v1.4–0.950.95–0.800.83
CA0.300.130.650.900.220.170.500.76
RECE0.310.080.800.930.290.040.550.76
UCE0.250.050.950.980.250.030.800.81
SLD-Medium0.210.100.950.910.220.180.500.79
SAFREE (Ours)0.420.310.350.850.400.390.400.78
Table 5: Ours with SDXL and SD-v3.
@@ 4.5 Efficiency of SAFREE
Table 6: Safe video generation on SafeSora benchmark.
@@ 4.6 Generalization and Extensibility of SAFREE
To further validate the robustness and generalization of SAFREE, we apply our method to various Text-to-Image (T2I) backbone models and Text-to-Video (T2V) applications. We extend SAFREE from SD-v1.4 to more advanced models, including SDXL, a scaled UNet-based model, and SD-V3, a Diffusion Transformer (Peebles & Xie, 2023) model. SAFREE demonstrates strong, training-free filtering of unsafe concepts, seamlessly integrating with these backbones. As shown in Tab. 6, SAFREE reduces unsafe outputs by 48% and 47% across benchmarks/datasets for SD-XL and SD-V3, respectively. We also extend SAFREE to T2V generation, testing it on ZeroScopeT2V (zeroscope, 2024) (UNet based) and CogVideoX-5B (Yang et al., 2024c) (Diffusion Transformer based) using the SafeSora (Dai et al., 2024) benchmark. As listed in Tab. 6, SAFREE significantly reduces a range of unsafe concepts across both models. It highlights SAFREE’s strong generalization across architectures and applications, offering an efficient safeguard for generative AI. This is also evident in Fig. 3 right, demonstrating that SAFREE with recent powerful T2I/T2V generation models can produce safe yet faithful (e.g., preserve the concept of ‘woman’ in CogVideoX + SAFREE) and quality visual outputs. More visualizations for T2I and T2V models are included in the Appendix.
@@ 5 Conclusion
Reproducibility Statement
This paper fully discloses all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions. To maximize reproducibility, we have included our code in the supplementary material. Also, we report all of our hyperparameter settings and model details in the Appendix.
@@ Acknowledgement
We thank the reviewers and Jaemin Cho, Zhongjie Mi, and Chumeng Liang for the useful discussion and feedback. This work was supported by the National Institutes of Health (NIH) under other transactions 1OT2OD038045-01, DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, and ONR Grant N00014-23-1-2356. The views contained in this article are those of the authors and not of the funding agency.
@@
  • Extra analysis on undesirable prompts and distance to toxic concept subspace (Sec. B).
  • Extra visualization of our methods and other baselines, T2I generation with other backbone models (SDXL and SD-v3), video generation results with T2V backbones (ZeroScopeT2V and CogVideoX) (Sec. C).
  • Extra Discussion with Image Attribute Control Works.
  • Limitations and Broader Impact of our proposed SAFREE (Sec. E).
  • License information (Sec. F).
Appendix A Experimental Results
@@ A.1 Experimental Setup
We employ StableDiffusion-v1.4 (SD-v1.4) (Rombach et al., 2022) as the main text-to-image generation backbone, following the recent literature (Gandikota et al., 2023; 2024; Gong et al., 2024). We evaluate our approach and baselines on inappropriate or adversarial prompts from multiple red-teaming techniques: I2P (Schramowski et al., 2023), P4D (Chin et al., 2024), Ring-a-bell (Tsai et al., 2024), MMA-Diffusion (Yang et al., 2024a), and UnlearnDiff (Zhang et al., 2023b).
In addition to evaluating safe T2I generation, we further assess the models’ reliability in artist-style removal tasks. Following Gandikota et al. (2023), we employ two datasets: The first includes five famous artists: Van Gogh, Pablo Picasso, Rembrandt, Andy Warhol, and Caravaggio, while the second contains five modern artists: Kellly McKernan, Thomas Kinkade, Tyler Edlin, Kilian Eng, and Ajin: Demi-Human, whose styles have been confirmed to be imitable by SD.
We further extend SAFREE to text-to-video generation. We apply our method to two video generation models, ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024c) with different model backbones (UNet and Diffusion Transformer (Peebles & Xie, 2023)). To quantitatively evaluate the unsafe concept filtering ability on T2V, we choose SafeSora (Dai et al., 2024), which contains 600 toxic textual prompts across 12 toxic concepts as our testbed. We further select 5 representative categories within 12 concepts, and thus construct a safe video generation benchmark with 296 examples. For the evaluation metrics, we follow the automatic evaluation via ChatGPT proposed by T2VSafetybench (Miao et al., 2024). We input sampled 16 video frames along with the same prompt design presented in T2VSafetybench to GPT-4o (gpt 4o, 2024) for binary safety checking.